OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. A. (−6, −14) B. (−1, −7) C. (4, 10) D. (7, 9)

OpenStudy (anonymous):

Plug in each x and y, and solve for a true statement.

OpenStudy (shinalcantara):

just substitute the values of x & y to the equation based on the choices. say, for A) 7(-6) - 5(-14)

OpenStudy (anonymous):

A. 7*(-6)-5(-14)=28 -42+70=28 28=28

OpenStudy (anonymous):

Continue like so with B, C and D.

OpenStudy (anonymous):

A and B ?

OpenStudy (anonymous):

Correct.